OpenStudy (francis1):

Find the mean, median and mode for the following data: 3, 9, 11, 3, 10, 12 Choose the correct order for listing these measures of central tendency from least to greatest. mode, median, mean median, mode, mean mean, median, mode mode, mean, median

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

To find the mean you add up all the numbers then divide by how many numbers there are. The mode is the number which there are two of them, there can be two modes or even none. So in this case it is 3. the median is when you line up all the numbers least to greatest and cross one of the least off then the greatest then a least then greatest then get the middle number

OpenStudy (anonymous):

The mode is 3, the mean is 8, and median is 9.5

OpenStudy (anonymous):

first put them in numerical order.... 3,3,9,10,11,12 The median is the middle number....(9 + 10)/2 = 19/2 = 9 1/2 the mode is the number that appears the most .....3 The mean is the average....add all numbers then divide by the # of numbers.... 3 + 3 + 9 + 10 + 11 + 12 = 48. and since there are 6 numbers, you divide by 6......48/6 = 8 so we have ...from least to greatest... mode = 3 mean = 8 median = 9.5 ANSWER : the last one

OpenStudy (francis1):

Using a line of best fit, which y-value would most likely be predicted for an x-value of 14? 7 10 12 15

OpenStudy (anonymous):

Well the line of best fit is considered to be positive so that mean the x and y value will be quite high in accordance with one another. So, its either 12 or 15

OpenStudy (anonymous):

However it cannot be 12 because it is already present and has an x value of 10.5, so the answer is 15

OpenStudy (anonymous):

yes...I agree.....15

OpenStudy (francis1):

The number of specialty cakes Kim sold each month during the past year is shown below. She created a box-and-whisker plot to see how consistent her business has been. What percentage of time were Kim's monthly cake orders less than 6 cakes per month? 100% 75% 50% 25%